1. The gas inside a balloon has a volume of 15.0 L at a pressure of 2.0 atm Calculate the pressure of the gas if is volume is compressed to 10.0 L at the same temperature


2. A given amount of oxyçen gas has a volume of 25.0 L at a temperature of 37°C and a pressure of 10 atm Atwhat temperature would this gas occupy a volume of 22.0 L at a pressure of 10 atm?


3. The volume of a gas sample at 0 degrees C and 1.0 atm is 10.0 L How many moles of gas are contained in the sample?



4. What would be the pressure of 6.40 g oxygen gas in a vessel with a volume of 4.5 L at 20°C?



5. A sample of oxygen gas, which is saturated with water vapor, is kept in a 10-L vessel at 30°C and has a pressure of 758 Torr If the pressure of the water vapor at this temperature s 31.8 Torr, what would be the pressure of the


dry oxygen?

1
Expert's answer
2021-11-11T15:04:02-0500

1) P1 V1= P2V2

2×15= P2×10

P2= "\frac{2×15}{10}=3atm"


2) T1= 37+273.15=310.15

Pressure is constant

"\frac{V_1}{T_1}= \frac{V_2}{T_2}"


"\frac{25}{310.15}=\frac{22}{T_2}"


"T_2 =\frac{22×310.15}{25}= 272.9K"


3) At "0\u00b0C" and 1 atm, 1 mole of any gas occupies 22.4 litres

Number of moles of 10litre at stp

="\frac{10}{22.4}= 0.4464 \> moles"


4)Number of moles of oxygen

"= \frac{6.40}{32}=0.2"


"Gas\> constant \>R=0
.082057l.atm/K.Mol"

T=20+273.15= 293.15K


PV=nRT

P="\frac{0.2×0.082057×293.15}{4.5}"

"=1.069\>atm"


5) Pressure due to Oxygen = Total pressure - vapour pressure


Pressure of dry Oxygen "758-31.8=726.2 torr"